calculate the coherence time and coherence length of white light of wave length range of 3500 angstrom to 6500 angstrom
1
Expert's answer
2020-11-30T15:22:35-0500

As per the given question,

The range of the wavelengths of the lights are

"\lambda_1=3500 A^\circ = 3500\times 10^{-10}m"

"\lambda_2=6500 A^\circ = 6500\times 10^{-10}m"

We know that,

"T=\frac{\lambda}{c}"

Hence,

"\Rightarrow T_1=\frac{3500\times 10^{-10}}{3\times 10^8}sec =1166.67\times 10^{-18}s"

"\Rightarrow T_1 = 1.167\times 10^{-15}s"

"T_2=\frac{6500\times 10^{-10}}{3\times 10^8}sec =2.167\times10^{-15}s"

Hence coherence time "(\Delta T) = (2.167\times 10^{-15}-1.167\times 10^{-15})"

"=1\times 10^{-15}sec"

coherence length "(l)=c\times \Delta T"

"\Rightarrow l=3\times 10^{8}\times10^{-15}m"

"\Rightarrow l = 3\times 10^{-7}m"